OpenStudy (anonymous):

Find the distance (-5,8) (2,-5)

OpenStudy (anonymous):

So here we are basically finding a hypotenuse. So we have... a^2+b^2=c^2 We have a and b of the sides, a is the distance of x, b is the distance of y. So... a=7 b=13 So if we put that in to our theorem, we get this. 7^2+13^2=c^2 Make sense?

OpenStudy (lgbasallote):

good try @Headdesk ! unfortunately, pythagorean theorem is for triangles yes? this is the distance between two points.

OpenStudy (anonymous):

Actually if you look at we do... make sense now?